### v1

# Enhancing Agentic RL with Progressive Reward Shaping and Value-based Sampling Policy Optimization

Abstract
Large Language Models (LLMs) empowered with Tool-Integrated Reasoning (TIR) can iteratively plan, call external tools, and integrate returned information to solve complex, long-horizon reasoning tasks. Agentic Reinforcement Learning (Agentic RL) optimizes such models over full tool-interaction trajectories, but two key challenges hinder effectiveness: (1) Sparse, non-instructive rewards, such as binary 0–1 verifiable signals, provide limited guidance for intermediate steps and slow convergence; (2) Gradient degradation in Group Relative Policy Optimization (GRPO), where identical rewards within a rollout group yield zero advantage, reducing sample efficiency and destabilizing training.
To address these challenges, we propose two complementary techniques: Progressive Reward Shaping (PRS) and Value-based Sampling Policy Optimization (VSPO). PRS is a curriculum-inspired reward design that introduces dense, stage-wise feedback — encouraging models to first master parseable and properly formatted tool calls, then optimize for factual correctness and answer quality. We instantiate PRS for short-form QA (with a length-aware BLEU to fairly score concise answers) and long-form QA (with LLM-as-a-Judge scoring to prevent reward hacking). VSPO is an enhanced GRPO variant that replaces low-value samples with prompts selected by a task-value metric balancing difficulty and uncertainty, and applies value-smoothing clipping to stabilize gradient updates.
Experiments on multiple short-form and long-form QA benchmarks show that PRS consistently outperforms traditional binary rewards, and VSPO achieves superior stability, faster convergence, and higher final performance compared to PPO, GRPO, CISPO, and SFT-only baselines. Together, PRS and VSPO yield LLM-based TIR agents that generalize better across domains.
Introduction
Large language models (LLMs) have shown remarkable capabilities in complex reasoning and tool-augmented decision-making, enabling applications such as question answering, planning, and interactive agents (zhao2023survey; huang2024understanding). When deployed as Tool-Integrated Reasoning (TIR) agents, LLMs are equipped with external tools — e.g., search APIs or knowledge bases — and can iteratively plan, call tools, and integrate their outputs to produce final answers. This multi-step, interleaved reasoning-with-action process is particularly effective in long-horizon tasks such as multi-hop question answering or problem solving in dynamic environments.
Training high-performing LLM-based TIR agents goes beyond single-turn reinforcement learning (zhang2025landscape). Agentic reinforcement learning (Agentic RL) explicitly optimizes the LLM policy over full tool-interaction trajectories, with feedback signals derived from the correctness and quality of the intermediate and final outputs. However, in practice, Agentic RL for TIR faces two key challenges:
  • Sparse and non-instructive rewards: Widely used reinforcement learning with verifiable rewards (RLVR) (liu2025reinforcement) often employs binary 0-1 signals to indicate pass/fail based on the final answer. These sparse signals provide limited guidance for improving intermediate tool-use steps, offer no notion of partial progress, and can lead to inefficient exploration and unstable optimization.
  • Gradient degradation in GRPO: Group Relative Policy Optimization (GRPO) (guo2025deepseek) is a strong baseline for multi-sample policy improvement, but when all rollouts in a group receive identical rewards, their computed advantages become zero. This results in dampened gradient updates, low sample efficiency, and unstable training dynamics.
To address these issues, we propose two complementary techniques for training LLM-based TIR agents:
  • Progressive Reward Shaping (PRS): a curriculum-inspired reward design that augments sparse verifiable rewards with dense intermediate signals, enabling the model to first master essential capabilities — generating parseable and correctly formatted tool calls — before optimizing more challenging objectives such as factual correctness and answer quality. In short-form QA, PRS incorporates a length-aware BLEU to avoid unfair penalization of correct short answers; in long-form QA, it integrates LLM-as-a-Judge scoring to mitigate reward hacking.
  • Value-based Sampling Policy Optimization (VSPO): an improved variant of GRPO that replaces low-value (zero-advantage) samples with prompts selected according to a task-value metric balancing uncertainty and difficulty. VSPO further applies value-smoothing clipping to maintain stable gradient scales even when high-value prompts are repeatedly sampled.
Experiments on long-form and short-form QA benchmarks demonstrate that PRS consistently outperforms traditional binary reward schemes, and VSPO achieves more stable training and superior performance compared to PPO, GRPO, CISPO, and SFT-only baselines. Together, PRS and VSPO yield LLM-based TIR agents that converge faster, generalize better across domains.
Our contributions are summarized as follows:
  • We design PRS, a general reward shaping framework that integrates stage-wise progression and dense intermediate feedback, instantiated for both short- and long-form QA.
  • We develop VSPO, a GRPO variant with value-based sampling and value smoothing clipping, improving sample efficiency and training stability and performance.
  • We empirically validate PRS and VSPO across multiple QA benchmarks, showing consistent gains over strong baselines.
Figure 1: Overview of PRS (rogressive Reward Shaping ) and VSPO(alue-based Sampling Policy Optimization).
Related Work
Tool-Integrated Reasoning.
Tool-Integrated Reasoning (TIR) (lin2025understanding) augments LLMs by enabling them to interleave natural-language reasoning with external tool calls, such as code execution or search, thereby improving adaptability in complex tasks. Early TIR systems based on in-context learning or SFT (yao2022react; schick2023toolformer) achieve promising results but often generalize poorly to unseen tasks and new tool configurations. Recent advances employ Reinforcement Learning with Verifiable Rewards (RLVR) (liu2025reinforcement), allowing agents to learn robust tool-use strategies through exploration, outperforming purely imitative methods. RL frameworks such as ToolRL (qian2025toolrl), AutoTIR (wei2025autotir), and VTool-R1 (wu2025vtool) interleave symbolic computation with language reasoning in unified rollouts, balancing precise tool-mediated operations with flexible verbal inference.
Agentic RAG.
Agentic RAG (ravuru2024agentic) enhances traditional RAG frameworks by embedding autonomous agents that govern the retrieval-generation process, thereby enabling dynamic decision-making, iterative refinement, and adaptive orchestration. Agents leverage established patterns such as reflection (madaan2023self), planning (madaan2023self), tool use, and multi-agent collaboration (guo2024large) to manage retrieval strategies and integrate diverse information sources. Architectures range from single-agent systems to multi-agent and hierarchical designs, with some frameworks incorporating graph-based reasoning (Agent-G (leeagent)) for richer contextual integration. Compared to static RAG, Agentic RAG achieves greater flexibility, scalability, and contextual accuracy in domains such as healthcare, finance, and education (singh2024revolutionizing).
Training paradigm.
Post-pretraining optimization of LLMs encompasses Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). SFT leverages high-quality expert trajectories to align model behavior through imitation, but often exhibits limited generalization capability on out-of-distribution inputs or tasks demanding novel reasoning patterns not present in the training data. In contrast, RL facilitates active exploration by LLMs, enabling them to identify and reinforce superior behaviors that enhance generalization performance. RL methods for LLMs can be categorized as critic-based or critic-free. Critic-based approaches like PPO (schulman2017proximal) and VAPO (yue2025vapo) use a value function to provide fine-grained feedback during training. Critic-free methods like GRPO (shao2024deepseekmath) and CISPO (chen2025minimax) use only sequence-level rewards, offering greater computational efficiency and scalability.
Preliminary
Problem Formalization
In this section, we provide a formal defination of TIR (Tool-Integrated Reasoning).
Given a question $q$ with Related background information $h$ (e.g., historical dialogue records in dialogue system), the environment $\varepsilon$ provides access to a set of tools $\mathcal{T}=\{t_{0},t_{1},\dots t_{n}\}$, the language model $\mathcal{M}$ can proactively and iteratively interact with the environment by reasoning and calling specific tools in $\mathcal{T}$, obtaining the corresponding tool results from $\varepsilon$. This process repeats until $\mathcal{M}$ provides the final answer to $q$. This interaction process can be expressed as a multi-step trajectory $\mathbb{T}$.
The steps in $\mathbb{T}$ can be divided into two categories: intermediate tool calls and final answer. The intermediate tool call step $\tau_{k}$ can be defined as
$\tau_{k}=(r_{k},tc_{k},o_{k}),$ (1)
where $r_{k}$, $tc_{k}$ and $o_{k}$ denote the reasoning, tool call and returned observation respectively. The complete trajectory $\mathbb{T}$ is then defined as
$\mathbb{T}=\tau_{1}\cdot\tau_{2}\cdot\dots\tau_{n}\cdot\tau_{\text{final}},$ (2)
[…]
Agentic Reinforcement Learning
Based on Problem Formalization, the agentic RL training objective is:
$\max_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D},\mathbb{T}\sim\pi_{\theta}(\cdot\mid q;\mathcal{E})}\left[r_{\phi}(q,\mathbb{T})\right]\\ -\beta\mathbb{D}_{kl}\left[\pi_{\theta}(\mathbb{T}\mid q;\mathcal{E})\|\pi_{\mathrm{ref}}(\mathbb{T}\mid q;\mathcal{E})\right],$ (3)
where $\mathbb{T}$ denotes the trajectory generated by $\mathcal{M}$, $\pi_{\theta}$, and $\pi_{ref}$ denotes the policy model and reference model, respectively. $r_{\phi}$ and $D_{kl}$ denotes the returns and KL divergence respectively.
There are two well-established policy-gradient RL methods used in Agentic RL, Proximal Policy Optimization (PPO) (schulman2017proximal) and Group Relative Policy Optimization (GRPO) (guo2025deepseek).
For PPO, based on the form of 3, the $r_{\phi}(q,\mathbb{T})$ of PPO is
$r_{\phi}(q,\mathbb{T})$ $=\frac{1}{N}\sum_{i=1}^{N}\min\left(\frac{\pi_{\theta}\left(a_{t}^{(i)}\mid s_{t}\right)}{\pi_{\theta_{\text{old}}}\left(a_{t}^{(i)}\mid s_{t}\right)}A\left(s_{t},a_{t}^{(i)}\right),\right$.
$\quad\left.\operatorname{clip}\left(\frac{\pi_{\theta}\left(a_{t}^{(i)}\mid s_{t}\right)}{\pi_{\theta_{\text{old}}}\left(a_{t}^{(i)}\mid s_{t}\right)},1-\epsilon,1+\epsilon\right)A\left(s_{t},a_{t}^{(i)}\right)\right),$ (4)
where $a_{t}^{(i)}\sim\pi_{\theta_{\text{old }}}\left(a\mid s_{t}\right)$ is the i-th sampled token from the old policy $\pi_{\theta_{\text{old }}}$. $A_{t}$ is the estimated advantage given by
$A\left(s_{t},a_{t}\right)=\mathcal{R}\left(s_{t},a_{t}\right)-V\left(s_{t}\right),$ (5)
where $V_{\theta}(s)$ is the learned value function, which is derived from a critic network.
For GRPO, the $r_{\phi}(q,\mathbb{T})$ of GRPO is
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\ \{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x)}$ (6)
$\quad\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\right$.
$\quad\quad\left.\min\big(w_{i,t}(\theta)\hat{A}_{i,t},\ \operatorname{clip}\big(w_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\hat{A}_{i,t}\big)\right],$
$w_{i,t}(\theta)=\frac{\pi_{\theta}\left(y_{i,t}\mid x,y_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(y_{i,t}\mid x,y_{i,<t}\right)},$
$\hat{A}_{i,t}=\hat{A}_{i}=\frac{R\left(x,y_{i}\right)-\operatorname{mean}\big(\{R(x,y_{j})\}_{j=1}^{G}\big)}{\operatorname{std}\big(\{R(x,y_{j})\}_{j=1}^{G}\big)}.$
Here $x_{i}$ is the task (A.K.A prompt) and $y_{i}$ is the multi-step trajectory (A.K.A rollout) defined is equation 2. All tokens in $y_{i}$ share the same advantage as $\hat{A}_{i}$.
Reward Shaping
Reward shaping belongs to the category of reward design, which is a reinforcement learning (RL) technique that augments the original reward function with additional signals to accelerate and guide the agent’s learning process, while preserving the optimal policy(ibrahim2024comprehensive). In many real-world RL tasks, rewards are sparse, delayed, or noisy, making it difficult for agents to effectively explore the environment and converge to desirable behaviors. Reward shaping addresses these challenges by incorporating additional incentives or penalties that provide intermediate feedback, thereby encouraging exploration toward promising states and actions.
Methodology
Our methodology consists of two components: Progressive Reward Shaping and Value-based Sampling.
Progressive Reward Shaping (PRS)
We proposed progressive reward shaping (PRS) based on the following key insights: a effective reward function should exhibit (1) richer partial ordering and (2) instructive progression, both of which can significantly improve learning efficiency.
Richer partial order. A reward function with richer partial ordering—also described as being more discriminative or dense—can distinguish between rollouts of similar quality rather than assigning identical rewards. For example, the commonly used Reinforcement Learning with Verifiable Rewards (RLVR) (liu2025reinforcement) typically employs a binary reward of 0 and 1. Such sparse binary rewards provide limited learning signals: the model only receives ”correct” or ”incorrect” feedback, without finer-grained ”better” or ”worse” signals to guide improvement. Furthermore, for GRPO, this sparsity increases the likelihood of generating zero-advantage samples (discussed further in  VSPO (Value-based Sampling Policy Gradient)).
More instructive. Inspired by curriculum Learning (soviany2022curriculum), we design the reward function to start with simple, fundamental rewards and progressively transition into more challenging criteria. This staged approach encourages the model to first learn basic behaviors before advancing to more complex reasoning patterns, facilitating more stable and efficient learning.
Based on these principles, we present reward design framework in retrieval-augmented question-answering scenarios. Without loss of generality, this reward framework is applicable to diverse agentic RL training scenarios, and we provide a generalized formulation that can be adapted to other domains.
Short-Form QA
Short-Form QA refers to a type of QA where the answers are typically brief and mostly consist of concrete entities (e.g. time, place and name).
The standard reward function of short-form QA is binary 0-1 reward based on EM (Exact Match) (jin2025search):
$r_{\phi}(x,y)=\operatorname{EM}\left(a_{\text{pred }},a_{\text{gold }}\right)$ (7)
where $a_{\text{pred}}$ is the extracted final answer from response $y$ and $a_{\text{gold}}$ is the ground truth answer.
For TIR, model $\mathcal{M}$ must generate the correct intermediate tool call step $\tau_{k}$ of $\mathbb{T}$ before producing the final answer step $\tau_{\text{final}}$. We identify three progressive learning objectives:
  1. 1.
    Complete process: each tool call $tc_{k}$ in $\tau_{k}$ can be successfully parsed and executed to return observations $o_{k}$ and the final step $\tau_{\text{final}}$ can be parsed to extract the final answer $a$.
  2. 2.
    Complete format: the content generated at each step to be correctly placed in the corresponding tags (e.g. <reasoning></reasoning>)
  3. 3.
    Correct answer: based on $\bigcup_{i}o_{i}$, $\mathcal{M}$ generates the correct answer $a$.
We design corresponding reward components for each learning objective. We will first introduce each reward and then provide the complete PRS formulation.
Process reward $R_{process}$. The definition of $R_{process}$ is:
[…]
Process reward evaluates if $\mathcal{M}$ can generate parseable tool calls and final answers.
Format reward $R_{format}$. The defination of $R_{format}$ is:
$R_{format}=\\ \begin{cases}0&\text{others}\\ 0.1&\text{Complete label package}\end{cases}$ (9)
The format reward evaluates whether $\mathcal{M}$ generates outputs that conforms to the required structured format.
Answer reward $R_{a}$. Unlike the binary EM reward in Eq. (7), we design a denser reward based on BLEU  (papineni2002bleu), named short-form BLEU, which is more suitable for short-form QA.
BLEU is a widely-used metric for machine translation evaluation. It measures the n-gram overlap between a predicted translation and one or more reference texts. The definition of BLEU is
$\mathrm{BLEU}$ $=\mathrm{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (10)
[…]
where $N=4$, $p_{n}$ is the modified $n$-gram precision, $w_{n}$ is the weight for each $n$-gram (usually $1/N$), $c$ is the candidate length, $r$ is the reference length, and $\mathrm{BP}$ is the brevity penalty.
However, standard BLEU with $N=4$ is problematic for short-form QA. When the predicted answer length is less than 4 tokens, even an exact match results in BLEU <1<1 due to undefined higher-order $n$-gram precisions. This artificially penalizes short answers and biases the model toward generating longer responses.
To address this issue, we propose short-form BLEU, which dynamically adjusts the $n$-gram weights $w_{n}$ according to the prediction length $c$. Specifically, let
$\max_{n}=\min(N,c),$ (11)
then we re-define the weights as
$w_{n}=\begin{cases}\frac{1}{\max_{n}},&n\leq\max_{n},\\ 0,&n>\max_{n}\end{cases}$ (12)
so that only valid $n$-grams up to $\max\_n$ contribute equally to the final score. This ensures that an exact match between a short predicted answer and the reference can achieve $\mathrm{BLEU}=1$, avoiding unintended reward degradation.
Complete Formula of PRS. The defination of short-formal PRS is
$PRS_{short}=\\ \begin{cases}R_{process}+R_{format}&\text{ if }R_{process}<1\\ R_{process}+R_{format}+R_{a}&\text{others}\end{cases}$ (13)
Note that $R_{process}+R_{format}+R_{a}\geq R_{process}+R_{format}$, so PRS encourage models to first have the ability to conduct complete interactions, and then improve the quality of final answer.
Long-Form QA
Different from short-form QA, relying solely on n-gram based metrics as answer reward is insufficient for long-form responses, as models exploit these metrics through reward hacking. To overcome this issue, we add a reward of LLM-as-a-Judge $R_{judge}$ on top of $PRS_{short}$. $R_{judge}$ mainly evaluates whether the model’s answers are factually grounded and free from hallucinations. The defination of $PRS_{long}$ is
$\begin{split}&PRS_{short}=\\ &\begin{cases}R_{process}+R_{format},&\text{if }R_{process}<1,\\ R_{process}+R_{format}+R_{judge},&\text{if }R_{process}\geq 1,\\ R_{process}+R_{format}+R_{judge}+R_{a},&\begin{aligned} &\text{if }R_{process}\geq 1\\ &\text{and }R_{judge}\geq 1\end{aligned}\end{cases}\end{split}$ (14)
$PRS_{long}$ encourage models to first generate a reasonable answer without illusions, and then maximize alignment with the reference answer.
The general form of PRS
PRS is not limited to QA scenarios and can serve as a design framework for various agentic RL training tasks. While $R_{\text{process}}$ and $R_{\text{format}}$ are universal components in TIR training, the framework can be extended with task-specific reward components. Let $\sigma(\cdot)$ be the sigmoid function, then the general form of PRS would be
$\begin{split}R_{\mathrm{PRS}}=&R_{1}+\mathbb{I}\!\left(R_{1}\geq\epsilon_{1}\right)\cdot\sigma(R_{2})\\ &+\mathbb{I}\!\left(R_{1}\geq\epsilon_{1},R_{2}\geq\epsilon_{2}\right)\sigma(\cdot R_{3})+\cdots.\end{split}$ (15)
Where $\epsilon_{i}$ is the reward threshold to control reaching the next reward step. The sigmoid function $\sigma(\cdot)$ ensures that the cumulative reward at each stage is strictly greater than at the previous stage (i.e., $\sigma(R_{i})\in(0,1)$), thereby encouraging the model to progressively learn increasingly complex behaviors while maintaining stable gradient signals.
VSPO (Value-based Sampling Policy Gradient)
GRPO suffers from the gradient degradation problem when all rollouts within their group have the same reward. The resulting advantages are zero, leading to zero policy gradients. This has two negative impacts.
  • the number of valid samples in a batch decreases, increasing the variance of the corresponding gradient estimation, which leads to unstable training.
  • the gradient signal of the batch is dampened, resulting in a decrease in sample efficiency.
To address this issue, we propose Value-based Sampling Policy Optimization (VSPO), an improved variant of GRPO. Unlike the dynamic sampling approach in DAPO (yu2025dapo), which may be computationally intensive. VSPO replaces zero advantage samples with selectively sampling in the current batch, maintaining computational efficiency while improving gradient quality.
VSPO
To ensure all samples in the batch produce effective gradients, VSPO first filters out the samples $x$ in the current batch with reward variance $Var\!\left(\{R(x,y_{i})\}_{i=1}^{G}\right)$ less than the threshold $\varepsilon$ (here we set $\varepsilon$ as $1\mathrm{e}{-}6$), then replaces them by resampling from the remaining samples.
Denote current batch as $B_{j}$, prompt $x_{i}=(h_{i},q_{i})\in B_{j}$ and the rollout group of $x$ is $G_{x_{i}}$. Let $\mu_{i}$ and $\sigma_{i}^{2}$ be the mean and variance of rewards within $G_{x_{i}}$. Then we have $B_{j}=B_{j}^{\sigma^{2}>=\varepsilon}\cup B_{j}^{\sigma^{2}<\varepsilon}$ and $B_{j}^{\sigma^{2}>=\varepsilon}\cap B_{j}^{\sigma^{2}<\varepsilon}=\emptyset$.
A naive approach would be uniform sampling from $B_{j}^{\sigma^{2}>=\varepsilon}$. However, we argue this can be improved. The insight of VSPO is ”prioritize samples with higher learning value”.
The learning value of task. During the training process of GRPO, there are two types of tasks are prone to zero advantages:
  • To easy task: all rollouts of the task achieve the same high reward, indicating the model $\mathcal{M}$ already proficient in solving this task.
  • To hard task: all rollouts receive the same low reward, which means model $\mathcal{M}$ lacks the necessary capability.
To measure the value of task $x$, we reinterpret the reward mean $\mu_{x}$ and varience $\sigma^{2}_{x}$ of rollouts group of $x$ in GRPO. $\sigma^{2}_{x}$ reflects the policy’s uncertainty in solving $x$. The higher the $\sigma^{2}_{x}$, the more gradient signals it contains. $\mu_{x}$ implies the average ability for $\mathcal{M}$ to solve $x$. The higher the $\mu_{x}$, the better $x$ is solved by $\mathcal{M}$.
Based on the above insights, we define the value score of $x$ as:
$V_{x}=(R_{max}-\mu_{x})\cdot\sigma^{2}_{x},$ (16)
where $R_{max}=\max_{x\in B,y_{i}\in G_{x}}R(x,y_{i})$ is the maximum reward of current batch $B$.
The term $(R_{\max}-\mu_{x})$ captures the difficulty of $x$ and represents the potential improvement gap between the best rollout and average performance, while $\sigma_{x}^{2}$ captures the policy’s uncertainty. This formulation balances both difficulty and uncertainty, ensuring that tasks with high learning potential receive appropriate sampling priority.
Value-based sampling Based on eaquation 16, we normalize the values using a temperature-controlled Softmax:
$p_{x_{i}}=\frac{\exp(V_{x_{i}}/T)}{\sum_{k}\exp(V_{x_{i}}/T)},\quad T>0,$ (17)
where smaller $T$ increases preference for high-score samples, while a larger $T$ yields more uniform probabilities.
Finally, we sample $\{x_{i}\}$ from $B^{\sigma^{2}>=\varepsilon}$ according to $\{p_{x_{i}}\}$ to replace the filtered samples.
Value smoothing clipping Value-based sampling alone can cause training instability. If prompt $x_{K}$ is sampled $N$ times, its contribution to the objective in Eq. (18) becomes:
$\mathcal{J}_{x_{K}}(\theta)$ $=N\!\times\!\sum_{\{y_{i}\}_{i=1}^{G}\!\sim\!\pi_{\theta_{\text{old}}}(\cdot\mid x_{K})}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}$
$\quad\times\sum_{t=1}^{|y_{i}|}\min\!\Big(w_{i,t}(\theta)\,\hat{A}_{i,t},$
$\quad\quad\operatorname{clip}\!\big(w_{i,t}(\theta),\,1-\epsilon,\,1+\epsilon\big)\hat{A}_{i,t}\Big)\Bigg]$
$=\sum_{\{y_{i}\}_{i=1}^{G}\!\sim\!\pi_{\theta_{\text{old}}}(\cdot\mid x_{K})}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}$
$\quad\times\sum_{t=1}^{|y_{i}|}\min\!\Big(w_{i,t}(\theta)\cdot N\cdot\hat{A}_{i,t},$
$\quad\quad\operatorname{clip}\!\big(w_{i,t}(\theta),\,1-\epsilon,\,1+\epsilon\big)\cdot N\cdot\hat{A}_{i,t}\Big)\Bigg]$ (18)
This shows that sampling $x_{K}$ NN times is equivalent to scaling its advantage $\hat{A}_{i}$ by $N$, which can destabilize policy gradients.
To stabilize training, we further proposed value smoothing clipping:
$A_{\text{new}}=\left(\alpha-\frac{\alpha-1}{N}\right)\cdot A,$ (19)
where $\alpha\in[1,+\infty]$ controls the the clipping magnitude (note that $\lim_{N\to+\infty}(\alpha-\frac{\alpha-1}{N})=\alpha$) When $N=1$, $A_{\text{new}}=A$; when $N>1$, AA is smoothly reduced to prevent over-dominance of repeated samples in gradient updates.
A Dynamic Advantage Weight Perspective of VSPO
VSPO can be interpreted as task-value-based reweighting rather than random sampling. Consider the general policy gradient (sutton1999policy) with advantage function $A_{t}$:
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\pi_{\theta}\left(y_{t}\mid y_{<t}\right)A_{t}\right].$ (20)
In GRPO, all tokens in the same trajectory share the same advantage, so:
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\pi_{\theta}\left(y_{t}\mid y_{<t}\right)\right]\cdot A_{x}.$ (21)
Let $\omega=\left(\alpha-\frac{\alpha-1}{N}\right)$, then
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\pi_{\theta}\left(y_{t}\mid y_{<t}\right)\right]\cdot\omega_{x}\cdot A_{x}.$ (22)
Thus, tasks with higher learning value receive larger advantage weights $\omega_{x}\cdot A_{x}$, effectively increasing the update step size for these tasks and amplifying policies with greater relative advantages.
Experimental Setup
To evaluate the effectiveness of PRS and VSPO in training LLM-based TIR agents, we conduct experiments on the following two types of long-horizon reasoning tasks: short-form QA and long-form QA. In these tasks, the agent has access to at least one information retrieval tool to answer the question.
In short-form QA, we use Qwen2.5-7B-Instruct and in long-form QA, we use the Qwen3-14B (qwen3technicalreport) as the training model. All training uses full parameter training.
We adopt Trinity-RFT (pan2025trinity) as our training framework.
For VSPO, we set $T=1$ and $\alpha=2$. The rollout num of VSPO, CISPO and GRPO is 5. The learning rate of all algorithms is 1e-6. Training is performed on a single node with 8 H100 GPUs. To optimize GPU memory usage, we enable gradient checkpointing and use Fully Sharded Data Parallel (FSDP) with CPU offloading.
We also use gradient checkpointing, FSDP offloading, and vLLM-based rollouts with the same hyperparameters as above. The rollout temperature and top-p values are both set to 1.0, and the KL divergence coefficient $\beta$ and clip ratio $\epsilon$ are fixed at 0.001 and 0.2.
Datasets
Long-Form QA.
We use proprietary data from our production system. We select conversation records from the online conversation database as RL rollout data. There are three types of queries in total:
  • $Q_{simple}$: A plain text query containing a single question.
  • $Q_{multim}$: A multimodal query containing both images and text.
  • $Q_{multiq}$: A query containing multiple sub-questions.
We manually annotate reference answers based on our internal knowledge base.
For expert data used in SFT, we use knowledge distillation from the Qwen3-235B-A22B (qwen3technicalreport) to construct high-quality trajectory data. Specifically, we collect online service logs from Qwen3-235B-A22B and apply rule-based filtering and LLM-as-a-Judge cleaning to produce expert demonstrations.
Short-Form QA.
Following (jin2025search), we use seven benchmark datasets: (1) General Question Answering: NQ , TriviaQA, and PopQA. (2) Multi-Hop Question Answering: HotpotQA, 2WikiMultiHopQA, Musique , and Bamboogle as Short-Form QA train and test data.
We merge the training sets of NQ and HotpotQA as our training set and evaluation is conducted on the test or validation sets of all seven datasets to assess both in-domain and out-of-domain performance.
Evaluation Metrics
Long-Form QA.
We use Qwen3-235B-A22B-Instruct-2507 (qwen3technicalreport) as an LLM judge to evaluate responses against manually annotated reference answers. The evaluation prompts are provided in Appendix B.
Short-Form QA.
We use Exact Match (EM) as the evaluation metric, computed against the gold answers provided by each benchmark dataset.
Baselines
To verify the effectiveness of VSPO, we compare it with GRPO, PPO (schulman2017equivalence), CISPO (chen2025minimax) and SFT-only. In long-form QA, We also report the performance of Qwen3-235B-A22B as a reference for strong LLM capabilities.
For the reward design, we compare PRS with rule-based 0-1 reward to demonstrate the effectiveness of PRS. For short-form QA, using exact match as rule-based 0-1 reward (this is SEARCH-R1).
Main Results
Long-Form QA
The main results comparing VSPO with baseline methods across are presented in Table 1. VSPO surpasses most RL baselines (GRPO, PPO, CISPO) and the SFT-only setting on $Q_{simple}$, $Q_{multiq}$ and $Q_{multim}$. Compared to untrained, $Q_{simple}$, $Q_{multiq}$ and $Q_{multim}$ achieve 18.26% and 81.24% and 10% average relative improvement, which demonstrating that VSPO yields more effective policy updates in long-horizon tool-use scenarios.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
Qwen3-235B-A22B 0.8125 0.600 0.500
untrained 0.6025 0.4 0.5
SFT-only 0.7 0.575 0.4
GRPO 0.700 0.575 0.475
PPO 0.6625 0.700 0.55
CISPO 0.6875 0.4875 0.4
VSPO 0.7125 0.725 0.55
Table 1: The performance results of VSPO with baselines in long-form QA.
Short-Form QA
Results are shown in Table2. We observe that: (1) Across all seven datasets, VSPO+PRS achieves the highest average score (0.419), outperforming all other RL baselines, including PPO+PRS and GRPO+PRS. This demonstrates the effectiveness of combining VSPO with the proposed progressive reward shaping in improving short-form reasoning tasks for LLM agents. (2) Comparing reward designs, PRS consistently improves performance over the rule-based exact match reward (EM) across PPO, GRPO, and VSPO, confirming that reward shaping can better guide long-horizon reasoning even in short-form question answering. (3) Overall, VSPO yields stable gains in most datasets without sacrificing performance on any, indicating strong robustness across both in-domain (NQ, HotpotQA) and out-of-domain (TriviaQA, PopQA, 2wiki, Musique and Bamboogle) benchmarks.
NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle Avg
Direct Inference 0.134 0.408 0.14 0.183 0.25 0.031 0.12 0.181
SFT 0.318 0.354 0.121 0.217 0.259 0.066 0.112 0.207
PPO+EM(SEARCH-R1) 0.393 0.610 0.397 0.370 0.414 0.146 0.368 0.385
GRPO+EM(SEARCH-R1) 0.429 0.623 0.427 0.386 0.346 0.162 0.400 0.396
VSPO+EM 0.433 0.623 0.425 0.396 0.350 0.162 0.390 0.397
PPO+PRS 0.410 0.610 0.400 0.386 0.410 0.157 0.400 0.396
GRPO+PRS 0.440 0.639 0.420 0.400 0.39 0.171 0.413 0.410
VSPO+PRS 0.440 0.645 0.416 0.408 0.410 0.171 0.435 0.419
Table 2: The performance results of VSPO and PRS with baselines in shor-form QA.
Analysis
We analyze some issues based on long-form QA, and the same applies to the case of short-form QA.
VSPO v.s. Other RL Algorithms
The training reward and eval reward of both RL algorithm are show in Firure 2.
Figure 2: Reward curves on the training and validation sets for RL algorithm in long-form QA.
VSPO achieved the best performance based on rewards on the training and validation sets. As the number of training steps increases, PPO gradually converges to the same level as VSPO and GRPO on training set.
Among all RL algorithms, VSPO was the first to receive the best validation reward, this indicates that filling in samples that cannot be helpful for learning through value-based sampling makes the model more focused on high-value samples, which helps to increase convergence speed and achieve better performance.
The Effective of PRS
To demonstrate the effectiveness of PRS, we use rule-baesd 0-1 reward as the baseline, using VSPO as the RL algorithm. This 0-1 reward use Qwen2.5-72B-Instruct as the verifier, determine whether the model output passes based on the annotated answers, if passed, the reward is 1, otherwise it is 0.
The training reward and eval reward of both reward design are show in Figure 3. Note that the scales of the two rewards are inconsistent, so we normalize PRS using the maximum and minimum values within the range of PRS. From curves on training set, we can see that, compared to the 0-1 reward, PRS achieved a relatively good convergence level earlier. And the reward of PRS on the validation set is also higher than 0-1 reward.
Figure 3: Reward curves on the training and validation sets for two reward designs.
The result on the test set is shown in Table 3. From this table we can see that PRS outperforms 0-1 reward on all query types, which proves the effectiveness of PRS.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
0-1 reward 0.6905 0.600 0.500
PRS 0.7125 0.725 0.550
Table 3: Results on the test set of 0-1 reward and PRS.
Figure 4 shows the entropy loss of VSPO when using 0-1 reward and PRS as training reward. Entropy loss reflects the uncertainty of policy model in selecting the current action. We observe that PRS reduces uncertainty to a lower level earlier than the binary reward and achieves better test set performance. This is not entropy collapse (which would manifest as degraded test metrics), but rather efficient convergence. Therefore, we infer that PRS, as a curriculum-learning-inspired reward design, effectively guides the policy’s exploration and enables faster convergence to high-quality strategies. In contrast, the binary reward leads to inefficient exploration under sustained high uncertainty.
Figure 4: Entropy loss of training using 0-1 reward and PRS.
Ablation Analysis of VSPO
To evaluate the effectiveness of value-based sample and value smoothing clipping of VSPO, we conduct experiments using no sample without clip (equivalent to GRPO), random sample without clip, value-based sample without clip, random sample with clip and value-based sample with clip (equivalent to VSPO), respectively. The result on the test set is shown in Table 4. The table demonstrates that removing either value-based sampling or value smoothing clipping significantly degrades performance, confirming the necessity of both components. Notably, any sampling strategy without clipping leads to training instability and poor performance, while value-based sampling with clipping achieves the best results across all query types.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
w/o sample, w/o clip
0.700 0.575 0.475
random sample, w/o clip
0.460 0.160 0.100
value-based sample, w/o clip
0.450 0.175 0.300
random sample, w/ clip
0.500 0.300 0.025
value-based sample, w/ clip
0.7125 0.725 0.550
Table 4: Results on the test set of different ablation settings.
We further analyze the impact of value-based sampling and value smoothing clipping through KL divergence, training rewards, and validation rewards. Figure 5 presents these metrics across different configurations. Two key observations emerge:
(1) Value smoothing clipping stabilizes training. Without clipping, both random sampling and value-based sampling exhibit abnormally high KL divergence, leading to extremely unstable training. In contrast, configurations with clipping maintain stable KL divergence throughout training, demonstrating that value smoothing clipping is essential for training stability.
(2) Value-based sampling is necessary for performance gains. From the training and validation reward curves, we observe that random sampling fails to improve performance even when combined with value smoothing clipping. However, value-based sampling with clipping achieves substantial improvements on both training and validation sets, confirming the necessity of intelligent sample selection based on learning value.
(a) KL Loss
(b) Reward on eval set
(c) Reward on train set
Figure 5: The KL Loss and Reward Curve.
Conclusions
In this work, we formalized the Tool-Integrated Reasoning (TIR) problem and proposed two complementary techniques to improve agentic RL in complex reasoning tasks: Progressive Reward Shaping (PRS) and Value-based Sampling Policy Optimization (VSPO).
PRS introduces a curriculum-based reward design that provides dense, stage-wise learning signals, enabling models to first master essential capabilities—such as generating parseable outputs and properly formatted responses—before optimizing for challenging objectives like answer quality and factual grounding. We instantiated PRS in both short-form and long-form QA, with the short-form version incorporating a length-aware BLEU to avoid unfair penalization of correct short answers, and the long-form version incorporating an LLM-as-a-Judge signal to address reward hacking in complex outputs.
To address gradient degradation in GRPO when rollouts have identical rewards, we developed VSPO, which uses value-based sample to prioritize updates on tasks with higher learning value and applies value smoothing clipping to stabilize policy gradients. This combination ensures that policy updates focus on the most informative samples while preventing destabilization from repeated high-value prompts.
Extensive experiments on multiple short-form and long-form QA benchmarks demonstrate that PRS consistently outperforms traditional binary rewards, and VSPO achieves more stable training, better sample efficiency and superior performance across policy optimization algorithms. Overall, this work provides a generalizable reward shaping and policy optimization in agentic RL with TIR, applicable beyond QA tasks to any multi-step reasoning scenarios.
Appendix A Prompt Template for QA Agent
Answer the given question. You must conduct reasoning inside <reasoning> and </reasoning> first every time you get new information. After reasoning, if you find you lack some knowledge, you can call wiki_search, and it will return the top searched results. You can search as many times as you want. If you find no further external knowledge needed, you can directly provide the answer inside <answer> and </answer> without detailed illustrations. For example, <answer> xxx </answer>.
Question: {query}
Appendix B Prompt Template for LLM-as-aJudge Evaluation
Role and Background
You are an experienced AI capability evaluation expert, with rigorous logic and keen observation skills, highly adept at assessing whether AI responses meet the requirements.
Now, an online travel platform has launched a Q&A Agent that provides answers to merchants questions.
Your Task
You will receive the following information:
[Merchants Question]: The question raised by the merchant.
[Q&A Agents Response]: The answer provided by the Q&A Agent.
[Reference Answer]: The standard answer corresponding to the merchants question.
You need to, strictly based on the [Evaluation Criteria], give an evaluation of the [Q&A Agents Response].
You must strictly follow the internal step-by-step thinking process, comprehensively analyze and evaluate the [Q&A Agents Response] based on the provided relevant information, and output the evaluation result in the specified format.
Your Internal Thinking Process (Please carry out internally, do not output)
Follow the steps below in your mind, analyzing step by step:
Analyze [Evaluation Criteria]
Carefully read the Evaluation Criteria, understand the core evaluation question and the requirements of each criterion.
Analyze [Merchants Question]
Analyze and understand the Merchants Question. Some questions may contain multiple sub-questions, others may include images (in URL form). If there are multiple questions, you need to understand the meaning of each.
Analyze [Q&A Agents Response]
Analyze and understand the Q&A Agents Response. If the Merchants Question contains multiple sub-questions, the Q&A Agents Response may contain multiple answers. You need to understand each answer, extract the information inside, determine which question it answers, and check if the model responded to all the questions.
Analyze [Reference Answer]
Analyze and understand the Reference Answer. If the Merchants Question contains multiple sub-questions, the Reference Answer will have answers to each question, and sometimes the order may differ from the order of questions in the Merchants Question. You need to first understand each answer, extract the information inside, match it with the corresponding question, and know the reference answer for each.
Perform Evaluation
Based on the above analysis results, use the [Evaluation Criteria] and [Reference Answer] to analyze and evaluate the [Q&A Agents Response].
Evaluation Criteria
Core evaluation question: Does the [Q&A Agents Response] match the [Reference Answer]?
Evaluation criteria:
Match: The Agents response must be semantically consistent with the Reference Answer. If the Merchants Question contains multiple sub-questions, the Q&A Agents Response must include an answer to each question, and each must be semantically consistent with the corresponding Reference Answer.
Mismatch: If the above "Match" criterion is not met, then it is judged as "Mismatch". That is, the Q&A Agents Response has significant deviations from the Reference Answer. For example, when the merchant asks multiple questions, the Q&A Agents Response only answers some of them, etc.
Output Format Requirement
You must strictly follow the format below when outputting, and directly provide the evaluation rating without any extra explanation or comments.
<Evaluation Rating>: